A total of $3800 is invested into two accounts paying 4% and 7% annual interest. In one year the combined interest on the accounts is $218. Find the amount invested at 7%.

Answer by mananth(16946)

Part I 4.00% per annum ------------- Amount invested =x
Part II 7.00% per annum ------------ Amount invested = y
3800
Interest----- 218.00

Part I 4.00% per annum ---x
Part II 7.00% per annum ---y
Total investment
x + 1 y= 3800 -------------1
Interest on both investments
4.00% x + 7.00% y= 218
Multiply by 100
4 x + 7 y= 21800.00 --------2
Multiply (1) by -4
we get
-4 x -4 y= -15200.00
Add this to (2)
0 x 3 y= 6600
divide by 3
y = 2200
Part I 4.00% $ 1600
Part II 7.00% $ 2200

CHECK
1600 --------- 4.00% ------- 64.00
2200 ------------- 7.00% ------- 154.00
Total -------------------- 218.00

m.ananth@hotmail.ca